A photon of light with a wavelength of 433 nm is emitted from an excited hydrogen atom. Through what transition did the electron in this atom undergo?
1
Expert's answer
2018-06-04T08:26:08-0400
1/λ=R(1/(n_i^2 )-1/(n_j^2 )).
(1/(n_i^2 )-1/(n_j^2 ))=1/λR=1/(433∙〖10〗^(-9) )(1.097∙〖10〗^7 ) =0.21
So,
1/2^2 -1/5^2 =0.21
Thus, it is 5 to 2 transition.